OpenStudy (anonymous):

Pistachios cost $6.50/pound and Almonds cost $8.00/pound. How much of each to make a 50lb mixture that costs $7.50/pound? Any help in the right direction would be great!

OpenStudy (anonymous):

This took a few to figure out. Bear with me while I show the work.

OpenStudy (anonymous):

Aright.

OpenStudy (anonymous):

Main problem setup: \[6.5x+8y=7.50(50)=375\] We know that x+y=50. Solving for y we get y=50-x. Substituting back into the original problem: \[6.5x+8(50-x)=375\] \[6.5x+400-8x=375\] \[-1.5x=-25\] \[x=16+2/3\] From the earlier equation we can calculate y: \[y=50-(16+2/3)=33+1/3\]

OpenStudy (anonymous):

If you plug the values back into the the original problem you get $375